prooe. During battery assembly, the electrodes are measured to the nearest tenth of a millimeter. The lengths are uniformly distributed with values at every tenth of a millimeter starting at 490.0 and continuing through 490.9. a) What is the mean of the lengths? b) What is variance of the lengths? 4.
Answer #1

This is an example of uniform distribution over (a,b)

where, a = 490.0 and b = 490.9

a) mean of the lengths

= mean of uniform distribution

= 1/2 * (a+b)

= 1/2 * (490.0+490.9)

= 490.45

b) variance of the lengths

= variance of uniform distribution

= 1/12 * (b-a)^2

= 1/12 * (490.9 - 490)^2

= 0.0675